OpenStudy (anonymous):

slope intercept of line containing (4, -1); parallel to 4x-9y=2

OpenStudy (agent47):

Slope intercept form is: y = mx+b, where m is the slope and b is the intercept

OpenStudy (anonymous):

i know. and parallel is the same slope. but whats really throwing me off is that its in standard form.

OpenStudy (agent47):

parallel to 4x-9y=2 means they have the same slope. To find it, get 4x-9y=2 into slope intercept form.

OpenStudy (agent47):

4x-9y=2 4x-9y-4x=2-4x -9y = -4x + 2 y = (4/9)x - (2/9) <- This is the form mx+b, where m = 4/9, thus the slope is 4/9.

OpenStudy (agent47):

So now you just use the following equation: y-y1=m(x-x1), where m is the slope, and (x1, y1) is the point.

OpenStudy (anonymous):

so its would be y=4/9(x-4)-1?

OpenStudy (agent47):

y - (-1) = (4/9)(x-4) y = (4/9)x-(16/9)-1 y = (4/9)x - (25/9) <- FINAL ANSWER

OpenStudy (anonymous):

could you continue to help me tonight? you seem to be the only nice person here...im not cheating. i just want to understand T_T